OpenStudy (anonymous):

Find the distance between the points given. Simplify all irrational answers. (0, 5) and (-5, 0)

OpenStudy (saifoo.khan):

What'd you got?

OpenStudy (anonymous):

show me step by step on how to do this... cause it says that the anwer is 

OpenStudy (saifoo.khan):

It's just same as that. what answer did you got? Just 1 more step.

OpenStudy (anonymous):

oh ok hold on let me work it

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

sqrt50

OpenStudy (saifoo.khan):

Correct. Now we will solve the radical. \[\sqrt{50} \to \sqrt{25 \times 2} \to \sqrt{5 \times 5\times 2}\to 5 \sqrt2\]

OpenStudy (anonymous):

ohhhhhh okay!!:D thanksssss!

OpenStudy (saifoo.khan):

Welcome. :)